eEgg is considering the purchase of a new distributed network computer system to help handle its warehouse inventories. The system costs $41,000 to purchase and install and $25,000 to operate each year. The system is estimated to be useful for 4 years. Management expects the new system to reduce the cost of managing inventories by $46,000 per year. The firm’s cost of capital (discount rate) is 11%.

Required:

1. What is the net present value (NPV) of the proposed investment under each of the following independent situations? (Use the appropriate present value factors from Appendix C, TABLE 1 and Appendix C, TABLE 2.)

1a. The firm is not yet profitable and therefore pays no income taxes.

1b. The firm is in the 21% income tax bracket and uses straight-line (SLN) depreciation with no salvage value. Assume MACRS rules do not apply.

1c. The firm is in the 21% income tax bracket and uses double-declining-balance (DDB) depreciation with no salvage value. Given a four-year life, the DDB depreciation rate is 50% (i.e., 2 × 25%). In year four, record depreciation expense as the net book value (NBV) of the asset at the start of the year.

2. What is the internal rate of return (IRR) of the proposed investment for situations in requirement 1, parts (a) through (c)? Use the IRR builit-in function in Excel to compute the IRR.

ESTIMATED IRR %
a
b
c
Answer #1
Ans 1a.
As there is no Tax, the depreciation can be ignored from cash flow as there will
be no Depreciation Tax shield.
NPV
Particulras Year 0 Year 1 Year 2 Year 3 Year 4
a Initial Investment              (41,000)
Cash flow from Operating Activities
Savings due to reduced Inventory Management cost                46,000              46,000            46,000             46,000
Less Operating Cost              (25,000)            (25,000)          (25,000)           (25,000)
b Total Cash flow from Operating activities                21,000              21,000            21,000             21,000
c Total Cash Flow =a+b=              (41,000)                21,000              21,000            21,000             21,000
d PV Factor @11%                         1               0.9009             0.8116           0.7312             0.6587
e PV of Cash flows=c*d=              (41,000)                18,919              17,044            15,355             13,833
f NPV =Sum of PV of Cash flows                24,150
g IRR (using excel function)= 36.44%
Ans 1 b.
NPV
Particulras Year 0 Year 1 Year 2 Year 3 Year 4
a Initial Investment              (41,000)
Cash flow from Operating Activities
Savings due to reduced Inventory Management cost                46,000              46,000            46,000             46,000
Less Operating Cost              (25,000)            (25,000)          (25,000)           (25,000)
Less Depreciation (SL)              (10,250)            (10,250)          (10,250)           (10,250)
Taxable Income                10,750              10,750            10,750             10,750
Less Tax @21%                  2,258                2,258              2,258               2,258
After Tax Income                  8,493                8,493              8,493               8,493
Add Back Depreciation                10,250              10,250            10,250             10,250
b Total Cash flow from Operating activities                18,743              18,743            18,743             18,743
c Total Cash Flow =a+b=              (41,000)                18,743              18,743            18,743             18,743
d PV Factor @11%                         1               0.9009             0.8116           0.7312             0.6587
e PV of Cash flows=c*d=              (41,000)                16,885              15,211            13,705             12,346
f NPV =Sum of PV of Cash flows                17,147
g IRR (using excel function)= 29.42%
Ans 1 c.
NPV Depreciation Caculation by DDB
Particulras Year 0 Year 1 Year 2 Year 3 Year 4 Year Opening Asset Book Value Depreciation @50% Closing Asset Book Value
a Initial Investment              (41,000) 1 $         41,000 $         20,500 $          20,500
Cash flow from Operating Activities 2 $         20,500 $         10,250 $          10,250